OpenStudy (anonymous):

find the domain and range f(x)=x+1/x-1

OpenStudy (tkhunny):

You have written \(f(x) = x + \dfrac{1}{x} - 1\). Is this what you intended?

OpenStudy (abb0t):

solve for the denominator, because you know that 1/0 isn't really true. In order to make it true, x must be greater than 1. x-1 = 0 x = 1

OpenStudy (anonymous):

@tkhunny no \[f(x) = x+1\div x-1\]

OpenStudy (anonymous):

@abb0t so what you mean the domain is: all real numbers except "1"? right?

OpenStudy (abb0t):

Yes. Correct.

OpenStudy (anonymous):

@abb0t if domain is D:{all real numbers\ 1} what is the range?

OpenStudy (anonymous):

\[f(x)=\frac{x+1}{x-1}\] domain all real numbers except \(x=1\) where the denominator is zero

OpenStudy (anonymous):

and what is the range? @satellite73

OpenStudy (anonymous):

range is also going to be all real number except 1, because a fraction is only equal to one if the numerator and denominator are the same, and here they cannot be equal

OpenStudy (anonymous):

thanksss

OpenStudy (anonymous):

yw

OpenStudy (tkhunny):

That's the same thing. Please make sure you know the difference between: x+1/x-1 and (x+1)/(x-1) Very different animals. The first is what you wrote. The second is what you intended.